# Code available at https://github.com/sulcantonin/neuro-symbolic-diagnostics.git

### v2

@@ Abstract
The development of intelligent agents, particularly those powered by language models (LMs), has shown the critical role in various environments that require intelligent and autonomous decision. Environments are not passive testing grounds and they represent the data required for agents to learn and exhibit very challenging conditions that require adaptive, complex and autonomous capacity to make decisions. While the paradigm of scaling models and datasets has led to remarkable emergent capabilities, we argue that scaling the structure, fidelity, and logical consistency of agent reasoning within these environments is a crucial, yet underexplored, dimension of AI research. This paper introduces a neuro-symbolic multi-agent architecture where the belief states of individual agents are formally represented as Kripke models. This foundational choice enables them to reason about known concepts of possibility and necessity using the formal language of modal logic. In this work, we use of immutable, domain-specific knowledge to make infere information, which is encoded as logical constraints essential for proper diagnosis. In the proposed model, we show constraints that actively guide the hypothesis generation of LMs, effectively preventing them from reaching physically or logically untenable conclusions. In a high-fidelity simulated particle accelerator environment, our system successfully diagnoses complex, cascading failures by combining the powerful semantic intuition of LMs with the rigorous, verifiable validation of modal logic and a factual world model and showcasing a viable path toward more robust, reliable, and verifiable autonomous agents.
@@ 1 Introduction
The trajectory of progress within the field of artificial intelligence has been influenced by the principles of scaling laws. The practice of scaling model size, dataset volume, and available computation has successfully allowed a wide range of emergent capabilities in large language models wei2022emergent . In a parallel intellectual movement, the field of agentic AI show that a similar scaling paradigm must be rigorously applied to the environments in which agents operate sapkota2025ai . The availability of rich, interactive, and diverse environments is foundational for the development of agents that can reason, and act with true autonomy. As the research focus shifts from agents trained on static datasets to those learning through sustained, dynamic environmental interaction via reinforcement learning (RL), the environment itself is transformed into the primary source for agent learning.
Simply making an environment look more realistic or adding more complexity is not enough to create the next generation of truly capable AI agents. For these agents to operate independently and reliably, especially in high-risk areas like industrial process control, automated scientific research and autonomous accelerator control sulc2024towards we need to understand how they think, be able to formally check their reasoning, and ensure their decision-making is firmly based on the core principles of the field they work in. Language Models (LMs) are powerful tools for processing and generating language, but they sometimes produce answers that sound convincing yet are factually wrong or logically flawed. This problem, often called hallucination, makes it risky to use them in critical systems where safety and reliability are essential ji2023survey ; zhang2025siren .
We propose a framework that addresses this pressing challenge by integrating two powerful computational paradigms: the creative hypothesis-generation capabilities of LMs and the strict, formal verification of symbolic logic. Specifically, we introduce a multi-agent architecture where each agent’s internal state is not a simple state vector but a formally structured Kripke model kripke1963semantical . This architectural choice allows agents to maintain and reason about a set of possible worlds, using the operators of modal logic $\Box$ for necessity and $\Diamond$ for possibility in order to formalize their beliefs about the environment’s current state and its potential future evolutions.
The core contributions of this work establish a comprehensive approach to building more reliable agents in high stake enviroments. First, we present a complete neuro-symbolic agent architecture where LMs are strategically employed for semantic interpretation and hypothesis generation, while modal logic provides the formal backbone for belief representation and state transition.
@@ 2.1 Kripke Models as Agent Belief States
First $W$ stands for a set of possible worlds. For a diagnostic agent, a ”world” represents a complete and plausible state of the entire environment. For instance, a world $w_{0}$ might represent a state where the system is operating nominally, while another world $w_{1}$ could represent a state where a specific cooling fault has occurred. Second element is $R$, which denotes an accessibility relation defined on $W$. A relation $(w_{i},w_{j})\in R$ signifies that if the agent currently believes it is in world $w_{i}$, it considers world $w_{j}$ to be a possible future or alternative state, thereby defining the agent’s uncertainty and its internal model of state transitions. Lastly, $V$ is a valuation function that assigns a set of true atomic propositions to each world $w\in W$. For example, the valuation of world $w_{1}$ might be $V(w_{1})=\{\text{pressure\_low},\text{cooling\_fault\_reported}\}$.
An agent’s current belief is always anchored to a specific world, $w_{\text{current}}\in W$. The agent then reasons about its beliefs by evaluating modal operators against this model. A proposition $p$ is considered to be possible, denoted $\Diamond p$, if there exists at least one world accessible from $w_{current}$ where $p$ is true. Conversely, a proposition is considered to be necessary, denoted $\Box p$, if $p$ is true in all worlds that are accessible from $w_{current}$. These logical rules are known to be decidable with a modal logic (without first order operators)
@@ 3 A Multi-Agent Architecture for Diagnostics
These serve as the observers, with each agent being responsible for a specific subsystem (e.g. cooling or radio-frequency power). These agents monitor a small, dedicated set of signals and utilize their local neuro-symbolic loop to form initial, localized hypotheses when anomalies are detected.
@@ 4 The Foundational Role of Expert Knowledge
This rule states that it is necessary that a cooling fault and a klystron fault are not the same event. In the context of the Kripke model, this means that for any world $w$ accessible from the current state, the valuation set $V(w)$ cannot simultaneously contain both propositions. This provides a logical boundary between different categories of failure.
@@ 6.1 Summary of Results
In Scenario 2 (Direct Causal Failure), the system demonstrated the power of its logical guardrails. The simultaneous faults triggered reports from both the ‘Klystron_Agent‘ and the ‘RF_Agent‘. The ‘AcceleratorDiagnostics‘ agent’s LM correctly proposed that the klystron was the root cause. This hypothesis was translated into a belief update that was tested against the axiom $\Box(\text{klystron\_fault\_reported}\rightarrow\text{rf\_power\_fault\_reported})$. The proposed theory was logically consistent with this rule and was therefore accepted and passed to the physical verification stage. LM incorrectly reversed the causality, the symbolic validation would have immediately failed because the resulting belief state would have violated this necessary implication. This successful test case shows the axioms functioning as intended to ensure physically sound reasoning.
@@ 7 Discussion and Future Work
Our work demonstrates a promising and viable path for building more reliable and trustworthy autonomous agents for a fast diagnosis at accelerator controls. By treating the LM not as an infallible oracle but as a powerful hypothesis generator whose outputs are subject to rigorous symbolic verification, we can effectively harness its strengths while simultaneously mitigating its known weaknesses. The use of modal logic and Kripke models provides a formal semantics for agent belief, which is a critical and necessary step towards creating agents that can transparently explain their reasoning and provide formal guarantees about their behavior. Further exploration into Dynamic Epistemic Logic could enhance the model’s ability to reason about how agents’ knowledge changes as a result of new information or actions van2007dynamic .
This framework opens up several exciting directions for future research. The current process of translating LM outputs into formal propositions is straightforward; developing more complex semantic parsing techniques, possibly using methodologies from the field of computational linguistics, could allow for the expression of richer and more nuanced hypotheses. Future work could explore advanced methods for learning these logical constraints directly from data or through guided interaction with human experts, potentially leading to systems that can formalize their own operational theories over time, a concept related to theory of mind in AI runge2020discovering . Finally, applying this architecture to an online reinforcement learning setting, where an agent’s actions have tangible consequences and its belief models are continually updated based on rich environmental feedback, represents the next logical step toward achieving true end-to-end autonomy.
@@ 8 Conclusion
As we work toward creating more autonomous and versatile AI agents, it’s clear they need environments that are scalable, interactive, and diverse. But scaling shouldn’t only mean bigger or more complex worlds — it must also include improving the agents’ ability to reason. For an agent to be truly reliable, it needs to build and maintain a logically consistent understanding of its environment. In this paper, we present a neuro-symbolic architecture that combines the language skills of language models (LMs) with the precision of modal logic and expert knowledge. This approach provides a framework that is both scalable and verifiable. By allowing agents to reason not just about what is true, but also about what is possible, necessary, or impossible, we take a significant step toward the long-standing goal of creating autonomous systems we can genuinely trust.
@@ A.2 Simplifications and Justifications
  • Linear and Discrete Processes: The physical relationships are simplified to linear functions, and the simulation progresses in discrete time steps (”ticks”). For instance, the temperature rise from a cooling fault is linear over time. This simplification ensures that the causal links are clear and unambiguous for the purpose of evaluating the diagnostic logic.

### v3

@@ Abstract
The development of intelligent agents, particularly those powered by language models (LMs), has shown a critical role in various environments that require intelligent and autonomous decision-making. Environments are not passive testing grounds, and they represent the data required for agents to learn and exhibit in very challenging conditions that require adaptive, complex, and autonomous capacity to make decisions. While the paradigm of scaling models and datasets has led to remarkable emergent capabilities, we argue that scaling the structure, fidelity, and logical consistency of agent reasoning within these environments is a crucial, yet underexplored, dimension of AI research. This paper introduces a neuro-symbolic multi-agent architecture where the belief states of individual agents are formally represented as Kripke models. This foundational choice enables them to reason about known concepts of possibility and necessity using the formal language of modal logic. In this work, we use immutable, domain-specific knowledge to make an informed root cause diagnosis, which is encoded as logical constraints essential for proper, reliable, and explainable diagnosis. In the proposed model, we show constraints that actively guide the hypothesis generation of LMs, effectively preventing them from reaching physically or logically untenable conclusions. In a high-fidelity simulated particle accelerator environment, our system successfully diagnoses complex, cascading failures by combining the powerful semantic intuition of LMs with the rigorous, verifiable validation of modal logic and a factual world model and showcasing a viable path toward more robust, reliable, and verifiable autonomous agents.
@@ 1 Introduction
The trajectory of progress within the field of artificial intelligence has been influenced by the principles of scaling laws. The practice of scaling model size, dataset volume, and available computation has successfully allowed a wide range of emergent capabilities in large language models wei2022emergent . In a parallel intellectual movement, the field of agentic AI shows that a similar scaling paradigm must be rigorously applied to the environments in which agents operate sapkota2025ai . The availability of rich, interactive, and diverse environments is foundational for the development of agents that can reason and act with true autonomy. As the research focus shifts from agents trained on static datasets to those learning through sustained, dynamic environmental interaction via reinforcement learning (RL), the environment itself is transformed into the primary source for agent learning.
Simply making an environment look more realistic or adding more complexity is not enough to create the next generation of truly capable AI agents. For these agents to operate independently and reliably, especially in high-risk areas like industrial process control, automated scientific research, and autonomous accelerator control sulc2024towards , we need to understand how they think, be able to formally check their reasoning, and ensure their decision-making is firmly based on the core principles of the field they work in. Language Models (LMs) are powerful tools for processing and generating language, but they sometimes produce answers that sound convincing yet are factually wrong or logically flawed. This problem, often called hallucination, makes it risky to use them in critical systems where safety and reliability are essential ji2023survey ; zhang2025siren .
We propose a framework that addresses this pressing challenge by integrating two powerful computational paradigms: the creative hypothesis-generation capabilities of LMs and the strict, formal verification of symbolic logic. Specifically, we introduce a multi-agent architecture where each agent’s internal state is not a simple state vector but a formally structured Kripke model kripke1963semantical . This architectural choice allows agents to maintain and reason about a set of possible worlds, using the operators of modal logic $\Box$ for necessity and $\Diamond$ for possibility to formalize their beliefs about the environment’s current state and its potential future evolutions.
The core contributions of this work establish a comprehensive approach to building more reliable agents in high-stakes environments. First, we present a complete neuro-symbolic agent architecture where LMs are strategically employed for semantic interpretation and hypothesis generation, while modal logic provides the formal backbone for belief representation and state transition.
@@ 2.1 Kripke Models as Agent Belief States
First, $W$ stands for a set of possible worlds. For a diagnostic agent, a ”world” represents a complete and plausible state of the entire environment. For instance, a world $w_{0}$ might represent a state where the system is operating nominally, while another world $w_{1}$ could represent a state where a specific cooling fault has occurred. The second element is $R$, which denotes an accessibility relation defined on $W$. A relation $(w_{i},w_{j})\in R$ signifies that if the agent currently believes it is in world $w_{i}$, it considers world $w_{j}$ to be a possible future or alternative state, thereby defining the agent’s uncertainty and its internal model of state transitions. Lastly, $V$ is a valuation function that assigns a set of true atomic propositions to each world $w\in W$. For example, the valuation of world $w_{1}$ might be $V(w_{1})=\{\text{pressure\_low},\text{cooling\_fault\_reported}\}$.
An agent’s current belief is always anchored to a specific world, $w_{\text{current}}\in W$. The agent then reasons about its beliefs by evaluating modal operators against this model. A proposition $p$ is considered to be possible, denoted $\Diamond p$, if there exists at least one world accessible from $w_{current}$ where $p$ is true. Conversely, a proposition is considered to be necessary, denoted $\Box p$, if $p$ is true in all worlds that are accessible from $w_{current}$. These logical rules are known to be decidable with a modal logic (without first-order operators)
@@ 3 A Multi-Agent Architecture for Diagnostics
These serve as the observers, with each agent being responsible for a specific subsystem (e.g., cooling or radio-frequency power). These agents monitor a small, dedicated set of signals and utilize their local neuro-symbolic loop to form initial, localized hypotheses when anomalies are detected.
@@ 4 The Foundational Role of Expert Knowledge
This rule states that a cooling fault and a klystron fault must not be the same event. In the context of the Kripke model, this means that for any world $w$ accessible from the current state, the valuation set $V(w)$ cannot simultaneously contain both propositions. This provides a logical boundary between different categories of failure.
@@ 6.1 Summary of Results
In Scenario 2 (Direct Causal Failure), the system demonstrated the power of its logical guardrails. The simultaneous faults triggered reports from both the ‘Klystron_Agent‘ and the ‘RF_Agent‘. The ‘AcceleratorDiagnostics‘ agent’s LM correctly proposed that the klystron was the root cause. This hypothesis was translated into a belief update that was tested against the axiom $\Box(\text{klystron\_fault\_reported}\rightarrow\text{rf\_power\_fault\_reported})$. The proposed theory was logically consistent with this rule and was therefore accepted and passed to the physical verification stage. LM incorrectly reversed the causality; the symbolic validation would have immediately failed because the resulting belief state would have violated this necessary implication. This successful test case shows the axioms functioning as intended to ensure physically sound reasoning.
@@ 7 Discussion and Future Work
Our work demonstrates a promising and viable path for building more reliable and trustworthy autonomous agents for fast diagnosis at accelerator controls. By treating the LM not as an infallible oracle but as a powerful hypothesis generator whose outputs are subject to rigorous symbolic verification, we can effectively harness its strengths while simultaneously mitigating its known weaknesses. The use of modal logic and Kripke models provides a formal semantics for agent belief, which is a critical and necessary step towards creating agents that can transparently explain their reasoning and provide formal guarantees about their behavior. Further exploration into Dynamic Epistemic Logic could enhance the model’s ability to reason about how agents’ knowledge changes as a result of new information or actions van2007dynamic .
This framework opens up several exciting directions for future research. The current process of translating LM outputs into formal propositions is straightforward; developing more complex semantic parsing techniques, possibly using methodologies from the field of computational linguistics, could allow for the expression of richer and more nuanced hypotheses. Future work could explore advanced methods for learning these logical constraints directly from data or through guided interaction with human experts, potentially leading to systems that can formalize their own operational theories over time runge2020discovering . Finally, applying this architecture to an online reinforcement learning setting, where an agent’s actions have tangible consequences and its belief models are continually updated based on rich environmental feedback, represents the next interesting step toward achieving true end-to-end autonomy.
@@ 8 Conclusion
As we work toward creating more autonomous and versatile AI agents, it’s clear that they need environments that are scalable, interactive, and diverse. But scaling shouldn’t only mean bigger or more complex worlds — it must also include improving the agents’ ability to reason. For an agent to be truly reliable, it needs to build and maintain a logically consistent understanding of its environment. In this paper, we present a neuro-symbolic architecture that combines the language skills of language models (LMs) with the precision of modal logic and expert knowledge. This approach provides a framework that is both scalable and verifiable. By allowing agents to reason not just about what is true, but also about what is possible, necessary, or impossible, we take a significant step toward the long-standing goal of creating autonomous systems we can genuinely trust.
@@ A.2 Simplifications and Justifications
  • Linear and Discrete Processes: The physical relationships are simplified to linear functions, and the simulation progresses in discrete time steps (”ticks”). For instance, the temperature rise from a cooling fault is linear over time. This simplification ensures that the causal links are unambiguous for the purpose of evaluating the diagnostic logic.